OpenStudy (anonymous):

Please Help! divide x^1/2 over x^2/7 my options are x^2/14 x^3/14 x^11/14 x^1/5

OpenStudy (saifoo.khan):

Use this,\[\Large \frac{x^a}{x^b} \to x^{a-b}\]

OpenStudy (anonymous):

(x^1/2)/(x^2/7) = x^(1/2 - 2/7)

OpenStudy (anonymous):

Thank you so much now I know how to find the answer!

OpenStudy (anonymous):

that's awesome